OpenStudy (anonymous):

One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side.

OpenStudy (anonymous):

square has sides of equal length so' 5n = n+16 5n-n=16 4n=16 n=16/4 n=4

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

the answer right?